PR3 v 280.17mV R1 PR2 PR1 11,266.82htv v-273.71mV 1002 100m 3 V1 C1 1V 10kHz 0° 4 10uF

Q: Calculate the amount of phase difference between the source voltage and the current in the circuit.
Answer #1

Total impedance of the circuit

Z=R+j( \omega L-\frac{1}{\omega C})

And the phase difference

\theta = tan^{-1} \frac{( \omega L-\frac{1}{\omega C})}{100}

\omega=2\pi f,\ f=10^4 Hz

After putting the value

\theta= 89.0879\degree